# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Type Ic Broad-Lined Supernova ?

### v1

@@ Abstract
Results. SN 2021lwz belongs to the rare class of rapidly evolving transients. The bolometric light curve rises in about days to a peak luminosity of about erg/s, at a rate of 0.2 mag day-1 close to the peak. Spectroscopy modeling reveals more similarities with a normal Type Ic-like SN than with a SLSN before peak, showing broadened lines after peak. Light curve modeling shows that the Arnett model of the bolometric light curve using a radioactive source (56 Ni) is not able to reasonably explain the light curve evolution. A magnetar model seems more appropriate, suggesting that the explosion of low ejecta mass () took place in a low mass () dwarf galaxy of specific star-formation rate about ten times larger than typical star-forming galaxies.
Conclusions. Given its spectroscopic properties and the low ejecta mass needed to model its light-curve, SN 2021lwz does not match with many core-collapse H-poor SNe Types. It shares similarities with rarer transients like SN 2014ft, iPTF 16asu and SN 2018gep.
@@
The advent of large-scale high-cadence surveys like Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), the Panoramic Survey Telescope and Rapid Response System (PanSTARRS; Kaiser et al. 2010), the Palomar Transient Factory (PTF) (see Rau et al. 2009), the Zwicky Transient Facility (ZTF; Bellm et al. 2019b; Graham et al. 2019) and the Asteroid Terrestrial- impact Last Alert System (ATLAS; Tonry et al. 2018a), developed over the last two decades, have opened the possibility to detect new types of transients. They have been useful to start populating the parameter space of the rising time, , defined as the time between the explosion and peak magnitude, versus absolute magnitude at peak, , with new populations of objects. Among those, superluminous supernovae (SLSNe) are explosions expected to originate from core collapsing stars. Such explosions may be forming a newborn rotating neutron star (Ostriker & Gunn 1971), or a magnetar (Kasen & Bildsten 2010; Woosley 2010). Interactions of the explosion products with the circumstellar medium (CSM) coming from previous progenitor mass loss is an additional mechanism often considered to interpret the diversity of light curves of H-poor SLSNe (e.g. Hosseinzadeh et al. 2022). The population of SLSNe is expected to fill the absolute magnitude range mag AB (see review by Gal-Yam 2019) or even a bit lower. For example, the light curves of the largest sample from ZTF (78 SLSNe) evolve slowly with a mean rest-frame rise time of 41.9 17.8 days (see Chen et al. 2023a, b). In this distribution low luminosity ( mag) extremely fast rising ( 10 days) transients are rare. From this sample, with a rising time of 10 days, and an absolute magnitude peaking close to , SN 2018bgv was the fastest SLSN observed by ZTF during the first months of the survey (Lunnan et al. 2020).
@@
The formation of stars as single objects, or in binary, and even triple stars systems, which is followed by their evolution up to a final stage (e.g. core collapse or disruptive events) suggest the existence of a complex diversity of star evolution channels. Stars evolving in binary systems are subject to strong interactions with their environment (e.g. binary star companion, black hole proximity, winds, shock waves, etc.) potentially leading to new types of transients. As such, a variety of new types of transients has recently been discovered among which Luminous Fast Blue Optical Transients (LFBOTs like AT 2018 (the Cow) and AT 2020xnd (the Camel); see for example Perley et al. 2021a; Ho et al. 2023, and references therein), Ultra-Stripped Supernovae (USSNe; like SN 2014ft (or iPTF 14gqr) and SN 2019ehk, see for example De et al. 2018; Nakaoka et al. 2021, and references therein), and Rapidly Evolving Transients (RETs, e.g. Drout et al. 2014; Arcavi et al. 2016; Tanaka et al. 2016; Pursiainen et al. 2018). Ideally, each of these types would be filling a given area in the phase space ( , ) but, in practice distinct populations sometimes fill similar areas leading to a degeneracy in this parameter space, and more data-driven information is required. This happens for example with well studied populations like Type Ia and Type II SNe, and with rarer populations like Type Ib, Ic and broad line Ic (or Ic-BL) SNe (see e.g. Figure 1 in Whitesides et al. 2017). Such a plot is presented in Figure 1 where the position of SN 2021lwz (see Section3.4) and several other objects of various types are displayed.
@@
The empirical method used to study such new populations of transients, and to try to understand the nature of their progenitors and their environments, is therefore to gather as much data as possible. This includes getting multi-band, multi-messenger, high cadence photometry allowing to estimate the full energy budget spatially deployed by the explosion over time (e.g. Prentice et al. 2016; Kornilov et al. 2023). This requires to get high cadence spectroscopy to study the chemical evolution of the photosphere (e.g. Quimby et al. 2011; Könyves-Tóth 2022; Könyves-Tóth & Seli 2023). When possible this means to get polarimetry data to study the time-evolution of the geometry of the systems closely connected to the explosion (e.g. inner core, jets, see Nagao et al. 2024; Soker 2024) or products resulting from the explosion (e.g. photosphere evolution, ejectae, see Inserra et al. 2016; Leloudas et al. 2017). Obviously gathering such data is a complex task even more complex for fast-rising events, and the lack of data obtained on some specific objects often prevents us from getting the full picture in one go. Nevertheless, gathering data and increasing the sample on several members of the same class of objects, retrospectively helps to better understand the said class. Such a progress is done in parallel to theoretical and numerical modeling developments (e.g. Khatami & Kasen 2024; Omand & Jerkstrand 2023), which are crucial to understand and interpret the data, and ultimately to put constraints on, or to identify the channel formation of such new or peculiar transients or SNe. In this context of luminous transients SN 2021lwz is a particularly intriguing object deserving a dedicated in depth study.
@@
SN 2021lwz (also known as ZTF21abaiono, PS21fcj, ATLAS21oii and Gaia21cvf), at R.A.09h44mn47.390s, Dec, was discovered by Fremling (2021) on May 10, 2021 from ZTF 111Zwicky Transient Facility, https://www.ztf.caltech.edu. public alerts (Patterson et al. 2019). The discovery magnitude obtained with the ZTF-cam mounted on the Palomar 1.2 meter Oschin was of g18.76 mag (AB system). The transient was classified on May 15, 2021 as a SLSN-I at a redshift z by Perley et al. (2021b) from the analysis of a Keck1 Low Resolution Imaging Spectrometer (LRIS, Oke et al. 1995) 222LRIS, https://www2.keck.hawaii.edu/inst/lris/lrishome.html spectrum. For that reason SN 2021lwz entered in our sample of H-poor SLSNe for follow-up. Subsequently, though, we realized that some of the pre-peak spectra are not similar to other typical H-poor SLSN spectra at similar phases, and we therefore decided to investigate this source in more detail. Indeed, SN 2021lwz is also part of the sample of rare Luminous SNe discussed by Gomez et al. (2022). This type of transient appearing to form a continuum between type Ic SNe and SLSNe is expected to make up of all observed Core Collapse SNe (CCSNe) (Gomez et al. 2022). Additionally, SN 2021lwz is a fast-rising SLSN ranging in the low absolute magnitude range, as do most of the He-rich SLSNe. This type is rare but a sample of seven Helium-Rich SLSNe was reported by Yan et al. (2020). In this sample, the spectroscopically well sampled SLSN SN 2019hge (ZTF 19aawfbtg) was shown to be of similar spectral type as PTF 10hgi, the only known He-rich SLSN at the time. Both were used to identify 6 more He-rich SLSNe thus increasing the sample of such known events to a total of 8 members.
@@
In this work we present an intensive study of SN 2021lwz. All the data obtained on SN 2021lwz and its host galaxy are presented in section 2. The data analysis of SN 2021lwz provided in section 3 is followed by detailed modeling of the SN and its host galaxy in section 4. In order to better understand and contextualize the nature of SN 2021lwz, comparisons with some SLSNe and other typical SNe and exotic transients are discussed in section 5. Our conclusions are given in section 6.
@@ 2 Data
In this section, we present data and information about the host galaxy. We also present the photometry, spectroscopy and polarimetry data obtained at several epochs on SN 2021lwz. Most of these observations were coordinated through the ZTF Fritz Marshal system (van der Walt et al. 2019; Coughlin et al. 2023). The reference epoch of maximum light is defined as MJD days with respect to g-band, as from a visual inspection. The extinction contribution from our Galaxy in the line-of-sight to SN 2021lwz and its host is and (Schlafly & Finkbeiner 2011).
@@ 2.1 Host Photometry
The highly accurate spectroscopic redshift z of SN 2021lwz was determined from narrow H-alpha host-galaxy emission and probable [OIII] emission at a redshift consistent with the less accurate redshift obtained with the SNID analysis (Perley et al. 2021b). The host galaxy at R.A.09h44mn47.397s, Dec is detected in archival Sloan Digitized Sky Survey (SDSS) 333http://skyserver.sdss.org/dr17/VisualTools/explore/summary and Legacy Survey (LS) 444Legacy Survey, https://www.legacysurvey.org/ images. It can be seen as the blue feature marked by the yellow circle in the LS image plotted in Figure 2.
We retrieved science-ready coadded images from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1) DR1 (Chambers et al. 2016) and the Sloan Digital Sky Survey data release 9 (SDSS DR 9; Ahn et al. 2012). In addition, the field was observed with the 3.58-m Canada France Hawaii Telescope telescope equipped with MegaPrime in u and r band between 2016 and 2019. We retrieved the level-2 data from the CADC and combined them using Swarp version 2.38.0 (Bertin 2010) to create a coadded image. We measure the brightness in all images using an aperture photometry tool presented in Schulze et al. (2018) that is based on Source Extractor (Bertin & Arnouts 2010). The measurements in these different bands are given in Table 4 (see Appendix A).
One survey observed that part of the sky in the near-infrared (NIR) with the Wide Field Camera (WFCAM) mounted on the United Kingdon Infrared Telescope 555UKIRT, https://about.ifa.hawaii.edu/ukirt/. The Large Area Scale (LAS) survey footprint of the UKIDSS Data Release 9 666https://cdsarc.cds.unistra.fr/viz-bin/cat/II/319 includes the coordinates position of SN 2021lwz. This survey whose limit magnitudes are all lower than 20.2 mag in the Y, J, H and K bands does not show any detection of the host of SN 2021lwz (see Lawrence et al. 2007), in agreement with our results from the modeling of the host (see Section 4.4).
@@ 2.2 Spectroscopy
In addition to the Keck1 LRIS public spectrum used to classify the transient, we acquired spectra with the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al. 2018) on the Palomar Observatory 60 inch telescope (P60), the LRIS on the 10 m Keck I telescope and the Andalucia Faint Object Spectograph and Camera (ALFOSC)777http://www.not.iac.es/instruments/alfosc/ on the 2.56 m Nordic Optical Telescope (NOT). SEDM spectra were reduced according to Rigault et al. (2019); Kim et al. (2022). The LRIS spectrum was reduced using LPipe as described by Perley (2019). ALFOSC spectra were reduced using PyNOT888https://github.com/jkrogager/PyNOT. A spectrophotometric standard star was used to calibrate each spectrum. A summary of the spectra obtained on SN 2021lwz is given in Table 5. In Figure 3 we show the spectral evolution of SN 2021lwz. The spectra were taken between rest frame days and days relative to the g-band maximum.
@@ 2.3 SN Photometry
Photometry of SN 2021lwz was obtained in several filters with the Swift satellite (Gehrels et al. 2004), with ZTF (Bellm et al. 2019b), with the Liverpool Telescope (LT) (Steele et al. 2004), and with the Asteroid Terrestrial-impact Last Alert System (ATLAS) 999https://atlas.fallingstar.com (Tonry et al. 2018b). Details about the data are given in the following sections. The photometry is shown in Figure 5. Table 6 summarises all these measurements.
@@ 3.1 Spectral analysis using template fitting
In order to get a first understanding of the spectral evolution of SN 2021lwz, we ran all the optical spectra through the SuperNova IDentifier (SNID) (Blondin & Tonry 2007, and references therein). We first performed a general classification by running the code over our upgraded template database which, in addition to the standard Types (Ia, Ib, Ic, II) and their subtypes, contains SLSN-I and SLSN-II templates including those of PTF 10hgi (19 spectra with epochs lying between -30 and +312 days with respect to peak in the SN rest-frame) and SN 2019hge (2 spectra at epochs 0 and +48 days, respectively) which are used as He-rich SLSN-I templates in Yan et al. (2020) (see Section 1).
The analysis provided by SNID shows that none of the spectra of SN 2021lwz obtained before peak (see Table 5 and Figure 3) can be matched properly without decreasing some of the values of the default SNID parameters aimed to ensure a statistically robust diagnostic. Fit attempts using only He-rich SLSNe PTF 10hgi or SN 2019hge templates were conducted on the pre-peak spectra without producing any good match. Using the full database, a relatively smooth fit is obtained on the -3.4 days Keck spectrum with the Type Ic-SL (Ic-superluminous) PTF10cwr, whose absorption features are smoothed compared to other SL templates. At maximum light the Keck spectrum at -0.5 days show similarities with the fast-evolving transient SN 2018bgv. These fits are shown in Figure 18.
Accurate fits of the ALFOSC spectra of SN 2021wz obtained post-peak are obtained without degrading the by-default SNID parameters. They suggest that the photosphere of SN 2021lwz evolved as a Ic-BL looking-like Type, in agreement with the first statement provided in the TNS classification report. We note that the fits obtained with PTF10cwr and SN 2018bgv templates shown in Figure 18; top, point to slightly lower redshifts than 0.065. This happens because SNID tries to adjust the W-feature often attributed to O ii, but it is known that different lines of other elements like Fe ii are expected in this wavelength range. This will be discussed further in our analysis. The fits obtained with SLSN-I SN 2016wi and with the Type Ic-SL SN LSQ12dlfd, shown in Figure 18; middle, point to more accurate redshifts consistency. Using a slightly lower value of 4 for the rlap factor, it is possible to produce lower quality fits with late time spectra of the peculiar Type Ic SN 1998bw (bottom) whose NIR spectral properties will be discussed and compared with those of SN 2021lwz NIR spectrum in Section 3.2.

### v2

# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Stripped Envelope Supernova ?

@@ Abstract
Results. SN 2021lwz belongs to the rare class of rapidly evolving transients. The bolometric light curve rises in about days to a peak luminosity of about erg/s, at a rate of 0.2 mag day-1 close to the peak. Spectroscopy modelling reveals more similarities with a normal Type Ic-like SN than with a SLSN before peak, showing slightly broadened lines after peak. Light curve modelling shows that the Arnett model of the bolometric light curve using a radioactive source (56 Ni) is not able to reasonably explain the light curve evolution. A magnetar model seems more appropriate, suggesting that the explosion of low ejecta mass () took place in a low mass () dwarf galaxy of specific star-formation rate about ten times larger than typical star-forming galaxies.
Conclusions. In conclusion SN 2021lwz is an uncommon transient showing many similarities with several classes of transients, and with rare transients. It may be an interesting example pointing on how differences in ejecta mass and engine parameters could produce a wide range of engine-driven SESNe.
@@
The advent of large-scale high-cadence surveys like Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), the Panoramic Survey Telescope and Rapid Response System (PanSTARRS; Kaiser et al. 2010), the Palomar Transient Factory (PTF) (see Rau et al. 2009), the Zwicky Transient Facility (ZTF; Bellm et al. 2019b; Graham et al. 2019) and the Asteroid Terrestrial- impact Last Alert System (ATLAS; Tonry et al. 2018a), developed over the last two decades, have opened the possibility to detect new types of transients. They have been useful to start populating the parameter space of the rising time, , defined as the time between the explosion and peak magnitude, versus absolute magnitude at peak, , with new populations of objects. Among those, superluminous supernovae (SLSNe) are explosions expected to originate from core collapsing stars. Such explosions may be forming a newborn rotating neutron star (Ostriker and Gunn 1971), or a magnetar (Kasen and Bildsten 2010; Woosley 2010). Interactions of the explosion products with the circumstellar medium (CSM) coming from previous progenitor mass loss is an additional mechanism often considered to interpret the diversity of light curves of H-poor SLSNe (e.g. Hosseinzadeh et al. 2022). A SLSN has often been considered as a type of stellar explosion that exhibits a peak luminosity ten or more times higher than that of a typical SN. This definition is quite general. In practice H-poor SLSNe are detected with absolute r-band magnitudes lying between and magnitudes (Gomez et al. 2024). The light curves of the largest sample from ZTF (78 SLSNe) evolve slowly with a mean rest-frame rise time of 41.9 17.8 days (see Chen et al. 2023a, b), while Gomez et al. (2024) derive a mean rise time of 27 days on a larger sample. In both distributions low luminosity ( mag) extremely fast rising ( 10 days) transients are rare. From the ZTF sample, with a rising time of 10 days, and an absolute magnitude peaking close to , SN 2018bgv was the fastest SLSN observed by ZTF during the first months of the survey (Lunnan et al. 2020).
@@
The formation of stars as single objects, or in binary, and even triple stars systems, which is followed by their evolution up to a final stage (e.g. core collapse or disruptive events) suggest the existence of a complex diversity of star evolution channels. Stars evolving in binary systems are subject to strong interactions with their environment (e.g. binary star companion, black hole proximity, winds, shock waves) potentially leading to new types of transients. As such, a variety of new types of transients has recently been discovered among which Luminous Fast Blue Optical Transients (LFBOTs like AT 2018cow (the Cow) and AT 2020xnd (the Camel); see for example Perley et al. 2021b; Ho et al. 2023, and references therein), Ultra-Stripped Supernovae (USSNe; like SN 2014ft (or iPTF 14gqr) and SN 2019ehk, see for example De et al. 2018; Nakaoka et al. 2021, and references therein). A more general terminology encompassing these events is Rapidly Evolving Transients (RETs, e.g. Drout et al. 2014; Arcavi et al. 2016; Tanaka et al. 2016; Pursiainen et al. 2018). Ideally, each of these types would be filling a given area in the phase space ( , ) but, in practice distinct populations sometimes fill similar areas leading to a degeneracy in this parameter space, and more data-driven information is required. This happens for example with well studied populations like Type Ia and Type II SNe, and with rarer populations like Type Ib, Ic and broad line Ic (or Ic-BL) SNe (see e.g. Figure 1 in Whitesides et al. 2017). Such a plot is presented in Figure 1 where the position of SN 2021lwz (see Section3.4) and several other objects of various types are displayed.
@@
The empirical method used to study such new populations of transients, and to try to understand the nature of their progenitors and their environments, is therefore to gather as much data as possible. This includes getting multi-band, multi-messenger, high cadence photometry allowing to estimate the full energy budget spatially deployed by the explosion over time (e.g. Prentice et al. 2016; Kornilov et al. 2023). This requires to get high cadence spectroscopy to study the chemical evolution of the photosphere (e.g. Quimby et al. 2011; Könyves-Tóth 2022; Könyves-Tóth and Seli 2023). When possible this means to get polarimetry data to study the time-evolution of the geometry of the systems closely connected to the explosion (e.g. inner core, jets, see Nagao et al. 2024; Soker 2024) or products resulting from the explosion (e.g. photosphere evolution, ejecta, see Inserra et al. 2016; Leloudas et al. 2017). Obviously gathering such data is a complex task even more complex for fast-rising events, and the lack of data obtained on some specific objects often prevents us from getting the full picture in one go. Nevertheless, gathering data and increasing the sample on several members of the same class of objects, retrospectively helps to better understand the said class. Such a progress is done in parallel to theoretical and numerical modelling developments (e.g. Khatami and Kasen 2024; Omand and Jerkstrand 2023), which are crucial to understand and interpret the data, and ultimately to put constraints on, or to identify the channel formation of such new or peculiar transients or SNe. In this context of luminous transients SN 2021lwz is a particularly intriguing object deserving a dedicated in depth study.
@@
SN 2021lwz (also known as ZTF21abaiono, PS21fcj, ATLAS21oii and Gaia21cvf), at R.A.09h44mn47.390s, Dec, was discovered by Fremling (2021) on May 10, 2021 from ZTF 111Zwicky Transient Facility, https://www.ztf.caltech.edu. public alerts (Patterson et al. 2019). The discovery magnitude obtained with the ZTF-cam mounted on the Palomar 1.2 meter Oschin was of g18.76 mag (AB system). The transient was classified on May 15, 2021 as a SLSN-I at a redshift z by Perley et al. (2021a) from the analysis of a Keck1 Low Resolution Imaging Spectrometer (LRIS, Oke et al. 1995) 222LRIS, https://www2.keck.hawaii.edu/inst/lris/lrishome.html spectrum. For that reason SN 2021lwz entered in our sample of H-poor SLSNe for follow-up. Subsequently, though, we realized that some of the pre-peak spectra are not similar to other typical H-poor SLSN spectra at similar phases, and we therefore decided to investigate this source in more detail. Indeed, SN 2021lwz is also part of the sample of rare Luminous SNe discussed by Gomez et al. (2022). This type of transient appearing to form a continuum between type Ic SNe and SLSNe is expected to make up of all observed Core Collapse SNe (CCSNe) (Gomez et al. 2022). Additionally, SN 2021lwz is a fast-rising SLSN ranging in the low absolute magnitude range, as do most of the He-rich SLSNe. This type is rare but a sample of seven Helium-Rich SLSNe was reported by Yan et al. (2020). In this sample, the spectroscopically well sampled SLSN SN 2019hge (ZTF 19aawfbtg) was shown to be of similar spectral type as PTF 10hgi, the only known He-rich SLSN at the time. Both were used to identify 6 more He-rich SLSNe thus increasing the sample of such known events to a total of 8 members.
@@
In this work we present an intensive study of SN 2021lwz. All the data obtained on SN 2021lwz and its host galaxy are presented in section 2. The data analysis of SN 2021lwz provided in section 3 is followed by detailed modelling of the SN and its host galaxy in section 4. In order to better understand and contextualize the nature of SN 2021lwz, comparisons with some SLSNe and other typical SNe and exotic transients are discussed in section 5. Our conclusions are given in section 6.
@@ 2 Data
In this section, we present data and information about the host galaxy. We also present the photometry, spectroscopy and polarimetry data obtained at several epochs on SN 2021lwz. Most of these observations were coordinated through the ZTF Fritz Marshal system (van der Walt et al. 2019; Coughlin et al. 2023). The reference epoch of maximum light is defined as MJD days with respect to g-band, as from a visual inspection. The extinction contribution from our Galaxy in the line-of-sight to SN 2021lwz and its host is and (Schlafly and Finkbeiner 2011).
@@ 2.1 Host photometry
The highly accurate spectroscopic redshift z of SN 2021lwz was determined from narrow H-alpha host-galaxy emission and probable [OIII] emission at a redshift consistent with the less accurate redshift obtained with the SNID analysis (Perley et al. 2021a). The host galaxy at R.A.09h44mn47.397s, Dec is detected in archival Sloan Digitized Sky Survey (SDSS) 333http://skyserver.sdss.org/dr17/VisualTools/explore/summary and Legacy Survey (LS) 444Legacy Survey, https://www.legacysurvey.org/ images. It can be seen as the blue feature marked by the yellow circle in the LS image plotted in Figure 2.
We retrieved science-ready co-added images from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1) DR1 (Chambers et al. 2016) and the Sloan Digital Sky Survey data release 9 (SDSS DR 9; Ahn et al. 2012). In addition, the field was observed with the 3.58-m Canada France Hawaii Telescope telescope equipped with MegaPrime in u and r band between 2016 and 2019. We retrieved the level-2 data from the CADC and combined them using Swarp version 2.38.0 (Bertin 2010) to create a co-added image. We measure the brightness in all images using an aperture photometry tool presented in Schulze et al. (2018) that is based on Source Extractor (Bertin and Arnouts 2010). The measurements in these different bands are given in Table 4 (see Appendix A).
One survey observed that part of the sky in the near-infrared (NIR) with the Wide Field Camera (WFCAM) mounted on the United Kingdom Infrared Telescope 555UKIRT, https://about.ifa.hawaii.edu/ukirt/. The Large Area Scale (LAS) survey footprint of the UKIDSS Data Release 9 666https://cdsarc.cds.unistra.fr/viz-bin/cat/II/319 includes the coordinates position of SN 2021lwz. This survey whose limit magnitudes are all lower than 20.2 mag in the Y, J, H and K bands does not show any detection of the host of SN 2021lwz (see Lawrence et al. 2007), in agreement with our results from the modelling of the host (see Section 4.4).
@@ 2.2 Spectroscopy
In addition to the Keck1 LRIS public spectrum used to classify the transient, we acquired spectra with the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al. 2018) on the Palomar Observatory 60 inch telescope (P60), the LRIS on the 10 m Keck I telescope and the Andalucia Faint Object Spectrograph and Camera (ALFOSC)777http://www.not.iac.es/instruments/alfosc/ on the 2.56 m Nordic Optical Telescope (NOT). SEDM spectra were reduced according to Rigault et al. (2019); Kim et al. (2022). The LRIS spectrum was reduced using LPipe as described by Perley (2019). ALFOSC spectra were reduced using PyNOT888https://github.com/jkrogager/PyNOT. A spectrophotometric standard star was used to calibrate each spectrum. A summary of the spectra obtained on SN 2021lwz is given in Table 5. In Figure 3 we show the spectral evolution of SN 2021lwz. The spectra were taken between rest frame days and days relative to the g-band maximum.
@@ 2.3 SN photometry
Photometry of SN 2021lwz was obtained in several filters with the Swift satellite (Gehrels et al. 2004), from the ZTF survey (Bellm et al. 2019b), and with the Liverpool Telescope (LT) (Steele et al. 2004). Public photometry was retrieved from the Asteroid Terrestrial-impact Last Alert System (ATLAS) 999https://atlas.fallingstar.com (Tonry et al. 2018b) public server (as explained in Section 2.3.4 ) Details about the data are given in the following sections. The photometry is shown in Figure 5. Table 6 summarises all these measurements.
@@ 3.1 Spectral analysis using template fitting
In order to get a first understanding of the spectral evolution of SN 2021lwz, we ran all the optical spectra through the SuperNova IDentifier (SNID) (Blondin and Tonry 2007, and references therein). We first performed a general classification by running the code over our upgraded template database which, in addition to the standard Types (Ia, Ib, Ic, II) and their subtypes, contains SLSN-I and SLSN-II templates including those of PTF 10hgi (19 spectra with epochs lying between -30 and +312 days with respect to peak in the SN rest-frame) and SN 2019hge (2 spectra at epochs 0 and +48 days, respectively) which are used as He-rich SLSN-I templates in Yan et al. (2020) (see Section 1).
The analysis provided by SNID shows that none of the spectra of SN 2021lwz obtained before peak (see Table 5 and Figure 3) can be matched properly without decreasing some of the values of the default SNID parameters aimed to ensure a statistically robust diagnostic. Fit attempts using only He-rich SLSNe PTF 10hgi or SN 2019hge templates were conducted on the pre-peak spectra without producing any good match. Using the full database, a relatively poor fit is obtained on the -3.4 days Keck spectrum with the Type Ic-SL (Ic-superluminous) PTF10cwr, whose absorption features are smoothed compared to other SL templates. At maximum light the Keck spectrum at -0.5 days show similarities with the fast-evolving transient SN 2018bgv. These fits are shown in Figure 19. Accurate fits of the ALFOSC spectra of SN 2021wz obtained post-peak are obtained without degrading the by-default SNID parameters. They suggest that the photosphere of SN 2021lwz evolved as a Ic-BL looking-like Type, in agreement with the first statement provided in the TNS classification report.
At early times, the spectra look (by eye) like they could be consistent with SLSNe, but with higher velocities than we are used to seeing. This could explain why the fits with SLSNe favour lower redshifts. This may also happen because SNID tries to adjust the W-feature often attributed to O ii, but it is known that different lines of other elements like Fe ii are expected in this wavelength range. This will be discussed further in our analysis. The fits obtained with SLSN-I SN 2016wi and with the Type Ic-SL SN LSQ12dlf, shown in Figure 19; middle, point to more accurate redshifts consistency. Using a slightly lower value of 4 for the rlap factor, it is possible to produce lower quality fits with late time spectra of the peculiar Type Ic SN 1998bw (bottom) whose NIR spectral properties will be discussed and compared with those of SN 2021lwz NIR spectrum in Section 3.2.